In 1993, Wildhorse Company completed the construction of a building at a cost of $2,360,000 and first occupied it in January 1994. It was estimated that the building will have a useful life of 40 years and a salvage value of $70,400 at the end of that time.

Early in 2004, an addition to the building was constructed at a cost of $590,000. At that time, it was estimated that the remaining life of the building would be, as originally estimated, an additional 30 years, and that the addition would have a life of 30 years and a salvage value of $23,600.

In 2022, it is determined that the probable life of the building and addition will extend to the end of 2053, or 20 years beyond the original estimate.

1. Using the straight-line method, compute the annual depreciation that would have been charged from 1994 through 2003. Annual depreciation from 1994 through 2003 :

2. Compute the annual depreciation that would have been charged from 2004 through 2022.Annual depreciation from 2004 through 2021:

3.Compute the annual depreciation to be charged, beginning with 2022. Annual depreciation expense—building :
Answer #1
1
Cost 2360000
Less: Salvage value 70400
Depreciable cost 2289600
Useful life 40
Annual depreciation from 1994 through 2003 57240 / yr.
2
Cost 2360000
Add: Additions 590000
Total cost 2950000
Less: Accumulated depreciation from 1994 through 2003 572400 =57240*10
Book value, Jan 2004 2377600
Less: Salvage value 94000 =70400+23600
Depreciable cost 2283600
Useful life 30
Annual depreciation from 2004 through 2021 76120 / yr.
3
Book value, Jan 2004 2377600
Less: Accumulated depreciation from 2004 through 2021 1370160 =76120*18
Book value, Jan 2022 1007440
Less: Salvage value 94000 =70400+23600
Depreciable cost 913440
Useful life 32 =12+20
Annual depreciation expense—building 28545